OpenStudy (anonymous):

3x divided by 7 - 2 = 15 what is x?

OpenStudy (anonymous):

x = 119/3

OpenStudy (anonymous):

25

jhonyy9 (jhonyy9):

3x/7 -2 =15 3x-14=105 3x=119 x=119/3